# El Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST data

### v1

@@
Gravitational lensing by galaxy clusters involves hundreds of galaxies over a large redshift range and increases the likelihood of rare phenomena (supernovae, microlensing, dark substructures, etc.). Characterizing the mass and light distributions of foreground and background objects often requires a combination of high-resolution data and advanced modeling techniques. We present the detailed analysis of El Anzuelo, a prominent quintuply imaged dusty star forming galaxy ($z_{\rm s}=2.29$), mainly lensed by three members of the massive galaxy cluster ACT-CL J0102--4915, also known as El Gordo ($z_{\rm d}=0.87$). We leverage JWST/NIRCam data containing previously unseen lensing features using a Bayesian, multi-wavelength, differentiable and GPU-accelerated modeling framework that combines Herculens (lens modeling) and NIFTy (field model and inference) software packages. For one of the deflectors, we complement lensing constraints with stellar kinematics measured from VLT/MUSE data. In our lens model, we explicitly include the mass distribution of the cluster, locally corrected by a constant shear field. We find that the two main deflectors (L1 and L2) have logarithmic mass density slopes steeper than isothermal, with $\gamma_{\rm L1}=2.23\pm 0.05$ and $\gamma_{\rm L2}=2.21\pm 0.04$. We argue that such steep density profiles can arise due to tidally truncated mass distributions, which we probe thanks to the cluster lensing boost and the strong asymmetry of the lensing configuration. Moreover, our three-dimensional source model captures most of the surface brightness of the lensed galaxy, revealing a clump of at most $400$ parsecs at the source redshift, visible at wavelengths $\lambda_{\rm rest}\gtrsim 0.6$ $\mu$m. Finally, we caution on using point-like features within extended arcs to constrain galaxy-scale lens models before securing them with extended arc modeling.
@@
A. ${}^{,}$ Corresponding author (aymeric.galan@gmail.com).El Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST data    G. B. Caminha El Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST data    J. Knollmüller El Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST data    J. Roth and S. H. Suyu El Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST data
@@ 1 Introduction
1.1 Strong gravitational lensing in galaxy clusters
Extending over megaparsec scales and reaching up to quadrillion solar masses, massive galaxy clusters are direct tracers of the cosmic web and its evolution through cosmic time. Galaxy clusters emerge as information-rich targets for jointly studying rare phenomena such as microlensing and caustic-crossing events (e.g., Dai et al. 2020; Williams et al. 2024), distant supernovae (e.g., Rodney et al. 2021; Frye et al. 2024), and probing the cosmological evolution of the Universe (e.g., Abbott et al. 2020; Bocquet et al. 2024). The most massive clusters are made of several hundreds of co-evolving galaxies, showcasing in a single scene the many different stages of galaxy evolution predominantly in the redshift range $z\sim 0-1$ (e.g., Wen et al. 2012; Hilton et al. 2021; Klein et al. 2023; Bulbul et al. 2024), probing the transition from a matter-dominated Universe to one driven by dark energy. Additionally, these clusters act as vast permeable screens standing between our telescopes and a plethora of more distant objects, enabling their detection up to redshifts $z\sim 13$ (e.g., Adams et al. 2023; Wang et al. 2023, 2024) and probing the epoch of reionization. Galaxy clusters do not only magnify these distant objects but also distort and duplicate their images, which can take the form of giant arcs extending over several arcseconds (Bayliss et al. 2011; Cava et al. 2018; Welch et al. 2022). This striking phenomenon, called strong gravitational lensing, is a direct consequence of the extensive gravitational potential of foreground galaxy clusters and their alignment with populations of background sources. A notable property of strong gravitational lensing in clusters is the extreme time delays between the multiple images of time-varying sources, which are used to geometrically measure absolute distances and cosmological parameters (e.g., Caminha et al. 2022; Kelly et al. 2023; Grillo et al. 2024), for which gains in precision are expected specifically with galaxy clusters (Acebron et al. 2023; Bergamini et al. 2024).
@@ 1.1 Strong gravitational lensing in galaxy clusters
1.2 The galaxy cluster El Gordo and the galaxy-scale strong lens El Anzuelo
@@
Significant improvements in lens models are achievable for a subset of cluster galaxies that coincidentally appear close in projection to extended lensed images of background sources, often visible as thick gravitational arcs. Alternatively, individual cluster members that are massive enough can create on their own additional strongly distorted and magnified extended arcs (so-called galaxy-galaxy strong lensing events in cluters, e.g., Meneghetti et al. 2023). These arcs are typically much fainter than the foreground lensing galaxies due to their larger distance, smaller intrinsic brightness and redshifted colors. Therefore, high signal-to-noise (S/N) and high resolution observations at wavelengths in which the lensed sources are brighter are necessary to provide useful constraints. In this regard, the James Webb Space Telescope (JWST) and its imaging capabilities in near-infrared wavelengths, have provided the clearest observations of massive galaxy clusters with resolved extended sources. One striking example is the famous merging galaxy cluster CT-CL J0102--4915 at redshift $z=0.87$, which has been one of the first clusters to be observed with JWST. Since its SZ-based discovery reported in Marriage et al. (2011), several independent analyses confirmed its total mass of $\gtrsim 10^{15}\ M_{\odot}$ giving it the nickname El Gordo as it is the most massive cluster at redshift $z\gtrsim 0.8$, when the Universe was approximately 6 Gyr old (Menanteau et al. 2012; Diego et al. 2023). Recent images acquired using the Near Infrared Camera (NIRCam) on JWST as part of the Prime Extragalactic Areas for Reionization and Lensing Science (PEARLS) Guaranteed Time Observing (GTO) program (Windhorst et al. 2023) have been publicly released in July 2023.
@@
The current state of mass measurements of El Gordo (using the SZ effect, weak lensing shape measurements, or strong lensing) is given in Table 1 of Diego et al. (2023). The first strong lensing analysis of the cluster have been achieved by Zitrin et al. (2013) with a light-traces-mass (LTM) model constrained by 9 families of multiple images deduced from Hubble Space Telescope (HST) imaging data. Taking advantage of deeper HST images and thus more image families, Diego et al. (2020) relaxed most of the modeling assumptions from previous works by performing a free-form model of the cluster, obtaining an independent mass measurement. More recently, Caminha et al. (2023) leveraged deep spectroscopic data obtained on the Very Large Telescope (VLT) with the Multi Unit Spectroscopic Explorer (MUSE) to measure the redshifts of 23 multiply lensed galaxies and 167 cluster members, significantly refining the lens model compared to previous works. In our analysis, we make use of both their cluster-scale model their VLT/MUSE data. Two subsequent modeling analysis of El Gordo have then been conducted using the JWST dataset from the PEARLS program. Diego et al. (2023) provided 28 new multiply imaged systems and obtained geometric redshifts (i.e., based on an initial lens model) for 37 systems, and then used the full set of 60 families to obtain a new free-form mass model. Shortly after, Frye et al. (2023) performed a new LTM model constrained by 56 multiply image systems to give further evidence of the two-component, cometary-like structure of the cluster, supporting that El Gordo is undergoing a major merger event.
@@
The JWST/NIRCam images of El Gordo revealed previously unseen features in lensed gravitational arcs, including two remarkable ones that have been recently studied in more details. One is the giant arc La Flaca that extends over 20 arseconds, making it one of the longest known gravitational arcs to date. Although already visible in previous HST images, JWST images contain numerous new multiply imaged clumps within the arc, providing constraints on intervening low-mass perturbers in the mass range $10^{9}-10^{10}\ M_{\odot}$ (Diego et al. 2023). Another remarkable lensing feature of El Gordo is called El Anzuelo, that we show in the top left panel of Fig. 1. El Anzuelo is a multiply imaged galaxy with a redshift of $z=2.29$ (based on CO line detection at millimeters wavelengths, Kamieneski et al. 2023) that appears as a highly distorted arc, bent around two cluster member galaxies (L1 and L2) and slightly influenced by a third one (L3). Previous HST images of El Anzuelo, covering mainly optical wavelengths, feature only one clear distorted image of the background source as shown in the top right panel of Fig. 1. The better sensitivity and resolution of JWST/NIRCam images in the near-infrared reveal that the source is, in fact, multiply imaged at least two times (we find that it is multiply imaged five times). The lensed source, also known as DSFG J010249--491507, is a dusty star forming galaxy whose multiple images have also been detected at radio wavelengths using the Atacama Large Millimeter/submillimeter Array (ALMA, see Cheng et al. 2023). In a recent work, Kamieneski et al. (2023) analyzed in detail the morphology and photometric properties of El Anzuelo using a combination of HST, JWST and ALMA datasets, and constructed a simple lens model based on compact multiply imaged features within the arcs. The authors measured the effective size of the lensed galaxy, characterized its dust distribution, and found that star formation is sensibly suppressed in its inner regions, supporting the hypothesis of ongoing inside-out quenching. We extensively compare our modeling results to the the analysis of Kamieneski et al. (2023).
@@
Figure 1: Color composite images of the dusty star forming galaxy El Anzuelo ($z_{\rm s}\approx 2.3$), strongly lensed by three members of the massive cluster El Gordo (L1, L2, L3, $z_{\rm d}\approx 0.9$). Top row, from left to right: color composite built from JWST/NIRCam data used in this work (red: F [^4] W; green: F410M, F356W; blue: F277W, F200W, F150W), archival HST/ACS+WFC3 data shown for comparison (red: F160W; green: F140W, F125W, F105W; blue: F775W, F625W, F606W, F435W). Bottom row, from left to right: best-fit model to the JWST/NIRCam data, unconvolved lensed source, unlensed source with arrows indicating the position of a clump revealed in our multi-band model.
1.3 This work: lens modeling of El Anzuelo using a 3D correlated field
@@
The main goal of our work is to conduct the first detailed analysis of the mass distribution of El Gordo members using El Anzuelo strong lensing features. To this end we should employ a model that is able to capture the complexity of the lensed galaxy over multiple wavelengths, in order to provide stronger constraints on the lens model. We propose to use Gaussian processes, which have been employed in various applications (both within and outside astrophysics) as a mathematically simple yet very versatile model to capture complex correlations observed in data, for an arbitrary number of dimensions. The specific implementation of Gaussian processes we consider here is developed in the Bayesian framework of Information Field Theory (IFT, Enßlin 2019) and is sometimes referred to as correlated fields. Some recent astrophysical applications of such fields include the 4D (spatial, time and frequency) reconstruction of the M87 supermassive black hole (Arras et al. 2022), the 3D map of line-of-sight galactic magnetic field in the Milky Way (Hutschenreuter et al. 2023), or the 3D map of the dust distribution in the solar neighborhood (Edenhofer et al. 2023). These recent works have demonstrated that Gaussian processes within IFT can be used to model complex correlations seen in various types of observational data (spatial, temporal, frequency). These results, as well as the efficiency of the accompanying algorithms (low computation times despite high model flexibility), motivate us to apply a similar strategy to model the highly detailed surface brightness of the lensed dusty star forming galaxy El Anzuelo.
@@
In our work we perform pixel-level and multi-wavelength modeling of the extended arcs of El Anzuelo as observed in the JWST/NIRCam dataset released by the PEARLS program. Building upon the work of Galan et al. (2022) that introduced the differentiable lens modeling code Herculens to describe complex mass and light distributions, we expand the code and use a 3D non-parametric correlated field model implemented in NIFTy (Selig et al. 2013; Steininger et al. 2019; Arras et al. 2019) to model the spatial and spectral morphology of El Anzuelo source galaxy. Prior to lens modeling, we use the software package starred (Michalewicz et al. 2023) to reconstruct the complex point spread function (PSF) of the instrument. Our modeling pipeline is Bayesian, differentiable and runs on GPU, which significantly improves the computation time necessary to model the many data pixels. We also complement our lensing constraints using stellar kinematics measured on recent MUSE/VLT spectroscopic data using the template fitting software package pPXF (Cappellari & Emsellem 2004; Cappellari 2017). To facilitate the access to our results and encourage future extensions of our analysis, we publicly release our models following the COOLEST lensing standard (Galan et al. 2023).
@@
Throughout this work, we assume a $\Lambda$CDM cosmology with $H_{0}=70\ {\rm km}\,{\rm s}^{-1}\,{\rm Mpc}^{-1}$ and $\Omega_{\rm m}=0.3$. This cosmology leads to angular sizes of approximately $7.7\ {\rm kpc\,arcsec}^{-1}$ at redshift $z=0.87$ and $8.2\ {\rm kpc\,arcsec}^{-1}$ at redshift $z=2.29$.

### v2

@@
A. Galan, Corresponding author (aymeric.galan@gmail.com).El Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST data    G. B. Caminha El Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST data    J. Knollmüller El Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST data    J. Roth and S. H. Suyu El Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST dataEl Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST data
Gravitational lensing by galaxy clusters involves hundreds of galaxies over a large redshift range and increases the likelihood of rare phenomena (supernovae, microlensing, dark substructures, etc.). Characterizing the mass and light distributions of foreground and background objects often requires a combination of high-resolution data and advanced modeling techniques. We present the detailed analysis of El Anzuelo, a prominent quintuply imaged dusty star-forming galaxy ($z_{\rm s}=2.29$), mainly lensed by three members of the massive galaxy cluster ACT-CL J0102--4915, also known as El Gordo ($z_{\rm d}=0.87$). We leverage JWST/NIRCam images, which contain lensing features that were unseen in previous HST images, using a Bayesian, multi-wavelength, differentiable and GPU-accelerated modeling framework that combines Herculens (lens modeling) and NIFTy (field model and inference) software packages. For one of the deflectors, we complement lensing constraints with stellar kinematics measured from VLT/MUSE data. In our lens model, we explicitly include the mass distribution of the cluster, locally corrected by a constant shear field. We find that the two main deflectors (L1 and L2) have logarithmic mass density slopes steeper than isothermal, with $\gamma_{\rm L1}=2.23\pm 0.05$ and $\gamma_{\rm L2}=2.21\pm 0.04$. We argue that such steep density profiles can arise due to tidally truncated mass distributions, which we probe thanks to the cluster lensing boost and the strong asymmetry of the lensing configuration. Moreover, our three-dimensional source model captures most of the surface brightness of the lensed galaxy, revealing a clump with a maximum diameter of $400$ parsecs at the source redshift, visible at wavelengths $\lambda_{\rm rest}\gtrsim 0.6$ $\mu$m. Finally, we caution on using point-like features within extended arcs to constrain galaxy-scale lens models before securing them with extended arc modeling.
@@ 1 Introduction
Extending over megaparsec scales and reaching up to quadrillion solar masses, massive galaxy clusters are direct tracers of the cosmic web and its evolution through cosmic time. Galaxy clusters emerge as information-rich targets for jointly studying rare phenomena such as microlensing and caustic-crossing events (e.g., Dai et al. 2020; Williams et al. 2024), distant supernovae (e.g., Rodney et al. 2021; Frye et al. 2024), and probing the cosmological evolution of the Universe (e.g., Abbott et al. 2020; Bocquet et al. 2024). The most massive clusters are made of several hundreds of co-evolving galaxies, showcasing in a single scene the many different stages of galaxy evolution predominantly in the redshift range $z\sim 0-1$ (e.g., Wen et al. 2012; Hilton et al. 2021; Klein et al. 2023; Bulbul et al. 2024), probing the transition from a matter-dominated Universe to one driven by dark energy. Additionally, these clusters act as vast permeable screens standing between our telescopes and a plethora of more distant objects, enabling their detection up to redshifts $z\sim 13$ (e.g., Adams et al. 2023; Wang et al. 2023, 2024) and probing the epoch of reionization. Galaxy clusters do not only magnify these distant objects but also distort and duplicate their images, which can take the form of giant arcs extending over several arcseconds (Bayliss et al. 2011; Cava et al. 2018; Welch et al. 2022). This striking phenomenon, called strong gravitational lensing, is a direct consequence of the extensive gravitational potential of foreground galaxy clusters and their alignment with populations of background sources. A notable property of strong gravitational lensing in clusters is the possible extreme time delays between the multiple images of time-varying sources (which can reach several years, see e.g., Li et al. 2012), which are used to geometrically measure absolute distances and cosmological parameters (see e.g., Caminha et al. 2022; Kelly et al. 2023; Grillo et al. 2024; Pascale et al. 2024), for which gains in precision are expected specifically with galaxy clusters (Acebron et al. 2023; Bergamini et al. 2024).
@@
In particular, extended arcs that form around individual galaxies can be used to probe the radial mass density slope $\gamma$ measured at the Einstein radius $\theta_{\rm E}$. Different processes can impact $\gamma$, such as the cooling of baryonic matter followed by dark matter halo contraction (e.g., Gnedin et al. 2004; Abadi et al. 2010), itself compensated for by outflows from stellar winds or AGN feedback (e.g., Johansson et al. 2012; Dubois et al. 2013). Large elliptical galaxies, which are prevalent of cluster-scale and galaxy-scale strong lenses, were found to have a radial behavior remarkably close to an isothermal profile (i.e., $\gamma=2$) from lensing and lens stellar dynamics constraints (e.g., Treu & Koopmans 2002; Auger et al. 2010; Sonnenfeld et al. 2013). This result has been termed the “bulge-halo conspiracy” in the literature (Treu et al. 2006; Dutton & Treu 2014; Etherington et al. 2023): despite the different radial behaviors of baryonic and dark matter profiles, their combination seems to lead to an isothermal density profile, with no evident signs for evolution with redshift (e.g., Sonnenfeld et al. 2013; Etherington et al. 2023; Sahu et al. 2024). At the population level, this result has been recently confirmed using lensing-only constraints, larger samples, and more modern lens modeling techniques (Shajib et al. 2021; Etherington et al. 2022; Tan et al. 2023).
@@
Significant improvements in lens models are achievable for a subset of cluster galaxies that coincidentally appear close in projection to extended lensed images of background sources, often visible as gravitational arcs. Alternatively, individual cluster members that are massive enough can create on their own additional strongly distorted and magnified extended arcs (so-called galaxy-galaxy strong lensing events in cluters, e.g., Meneghetti et al. 2023). These arcs are typically much fainter than the foreground lensing galaxies due to their larger distance, smaller intrinsic brightness and redshifted colors. Therefore, high signal-to-noise (S/N) and high resolution observations at wavelengths in which the lensed sources are brighter are necessary to provide useful constraints. In this regard, the James Webb Space Telescope (JWST) and its imaging capabilities in near-infrared wavelengths, have provided the clearest observations of massive galaxy clusters with resolved extended sources. One striking example is the famous merging galaxy cluster ACT-CL J0102--4915 at redshift $z=0.87$, which has been one of the first clusters to be observed with JWST. Since its SZ-based discovery reported in Marriage et al. (2011), several independent analyses confirmed its total mass of $\gtrsim 10^{15}\ M_{\odot}$ giving it the nickname El Gordo as it is the most massive known cluster at redshift $z\gtrsim 0.8$, when the Universe was approximately 6 Gyr old (Menanteau et al. 2012; Diego et al. 2023). Recent images acquired using the Near Infrared Camera (NIRCam) on JWST as part of the Prime Extragalactic Areas for Reionization and Lensing Science (PEARLS) Guaranteed Time Observing (GTO) program (Windhorst et al. 2023) have been publicly released in July 2023.
@@
The current state of El Gordo mass measurements (using the SZ effect, weak lensing shape measurements, or strong lensing) is given in Table 1 of Diego et al. (2023). The first strong lensing analysis of the cluster have been achieved by Zitrin et al. (2013) with a light-traces-mass (LTM) model constrained by 9 families of multiple images deduced from Hubble Space Telescope (HST) imaging data. Taking advantage of deeper HST images and thus more image families, Diego et al. (2020) relaxed most of the modeling assumptions from previous works by performing a free-form model of the cluster, obtaining an independent mass measurement. More recently, Caminha et al. (2023) leveraged deep spectroscopic data obtained on the Very Large Telescope (VLT) with the Multi Unit Spectroscopic Explorer (MUSE) to measure the redshifts of 23 multiply lensed galaxies and 167 cluster members, significantly refining the lens model compared to previous works. In our analysis, we make use of both their cluster-scale model their VLT/MUSE data. Two subsequent modeling analysis of El Gordo have then been conducted using the JWST dataset from the PEARLS program. Diego et al. (2023) provided 28 new multiply imaged systems and obtained geometric redshifts (i.e., based on an initial lens model) for 37 systems, and then used the full set of 60 families to obtain a new free-form mass model. Shortly after, Frye et al. (2023) performed a new LTM model constrained by 56 multiply image systems to give further evidence of the two-component, cometary-like structure of the cluster, supporting that El Gordo is undergoing a major merger event.
@@
The JWST/NIRCam images of El Gordo revealed features in lensed gravitational arcs that were unseen in previous HST images, including two remarkable ones that have been recently studied in more details. One is the giant arc La Flaca that extends over 20 arseconds, making it one of the longest known gravitational arcs to date. Although already visible in previous HST images, JWST images contain numerous new multiply imaged clumps within the arc, providing constraints on intervening low-mass perturbers in the mass range $10^{9}-10^{10}\ M_{\odot}$ (Diego et al. 2023). Another remarkable lensing feature of El Gordo is called El Anzuelo, that we show in the top left panel of Fig. 1. El Anzuelo is a multiply imaged galaxy with a redshift of $z=2.29$ (based on the most likely CO line detection at millimeters wavelengths, Kamieneski et al. 2023) that appears as a highly distorted arc, bent around two cluster member galaxies (L1 and L2) and slightly influenced by a third one (L3). Previous HST images of El Anzuelo, covering mainly optical wavelengths, feature only one clear distorted image of the background source as shown in the top right panel of Fig. 1. The better sensitivity and resolution of JWST/NIRCam images in the near-infrared reveal that the source is, in fact, multiply imaged at least two times (we find that it is multiply imaged five times). The lensed source, also known as DSFG J010249--491507, is a dusty star-forming galaxy whose multiple images have also been detected at radio wavelengths using the Atacama Large Millimeter/submillimeter Array (ALMA, see Cheng et al. 2023). In a recent work, Kamieneski et al. (2023) analyzed in detail the morphology and photometric properties of El Anzuelo using a combination of HST, JWST and ALMA datasets, and constructed a simple lens model based on compact multiply imaged features within the arcs. The authors measured the effective size of the lensed galaxy, characterized its dust distribution, and found that star formation is sensibly suppressed in its inner regions, supporting the hypothesis of ongoing inside-out quenching. We extensively compare our modeling results to the analysis of Kamieneski et al. (2023).
@@
Figure 1: Color composite images of the dusty star-forming galaxy El Anzuelo ($z_{\rm s}\approx 2.3$), strongly lensed by three members of the massive cluster El Gordo (L1, L2, L3, $z_{\rm d}\approx 0.9$). Top row, from left to right: color composite built from JWST/NIRCam data used in this work (red: F [^4] W; green: F410M, F356W; blue: F277W, F200W, F150W), archival HST/ACS+WFC3 data shown for comparison (red: F160W; green: F140W, F125W, F105W; blue: F775W, F625W, F606W, F435W). The exact weighting used for color composite images is given in Appendix A. Bottom row, from left to right: best-fit model to the JWST/NIRCam data, unconvolved lensed source, unlensed source with arrows indicating the position of a clump revealed in our multi-band model.
@@
The main goal of our work is to conduct the first detailed analysis of the mass distribution of individual El Gordo members using El Anzuelo strong lensing features. To this end we should employ a model that is able to capture the complexity of the lensed galaxy over multiple wavelengths, in order to provide stronger constraints on the lens model. We propose to use Gaussian processes, which have been employed in various applications (both within and outside astrophysics) as a mathematically simple yet very versatile model to capture complex correlations observed in data, for an arbitrary number of dimensions. The specific implementation of Gaussian processes we consider here is developed in the Bayesian framework of Information Field Theory (IFT, Enßlin 2019) and is sometimes referred to as “correlated fields”. Some recent astrophysical applications of such fields include the 4D (spatial, time and frequency) reconstruction of the M87 supermassive black hole (Arras et al. 2022), the 3D map of line-of-sight galactic magnetic field in the Milky Way (Hutschenreuter et al. 2023), or the 3D map of the dust distribution in the solar neighborhood (Edenhofer et al. 2024). These recent works have demonstrated that Gaussian processes within IFT can be used to model complex correlations seen in various types of observational data (spatial, temporal, frequency). These results, as well as the efficiency of the accompanying algorithms (low computation times despite high model flexibility), motivate us to apply a similar strategy to model the highly detailed surface brightness of the lensed dusty star-forming galaxy El Anzuelo.
@@
In our work, we perform pixel-level and multi-wavelength modeling of the extended arcs of El Anzuelo as observed in the JWST/NIRCam dataset released by the PEARLS program. Building upon the work of Galan et al. (2022) that introduced the differentiable lens modeling code Herculens to describe complex mass and light distributions, we expand the code and use a 3D non-parametric correlated field model implemented in NIFTy (Selig et al. 2013; Steininger et al. 2019; Arras et al. 2019) to model the spatial and spectral morphology of El Anzuelo source galaxy. Prior to lens modeling, we use the software package starred (Michalewicz et al. 2023) to reconstruct the complex point spread function (PSF) of the instrument. Our modeling pipeline is Bayesian, differentiable and runs on GPU, which significantly improves the computation time necessary to model the many data pixels. We also complement our lensing constraints using stellar kinematics measured on recent MUSE/VLT spectroscopic data using the template fitting software package pPXF (Cappellari & Emsellem 2004; Cappellari 2017). To facilitate the access to our results and encourage future extensions of our analysis, we publicly release our models following the COOLEST lensing standard (Galan et al. 2023).
@@
Throughout this work, we assume a Lambda cold dark matter ($\Lambda$CDM) cosmology with $H_{0}=70\ {\rm km}\,{\rm s}^{-1}\,{\rm Mpc}^{-1}$ and $\Omega_{\rm m}=0.3$. This cosmology leads to angular sizes of approximately $7.7\ {\rm kpc\,arcsec}^{-1}$ at redshift $z=0.87$ and $8.2\ {\rm kpc\,arcsec}^{-1}$ at redshift $z=2.29$.